OpenStudy (anonymous):

*Multiply and simplify* SQRT5 x SQRT 5

OpenStudy (anonymous):

SQRT5 x SQRT 5=5

sam (.sam.):

\[\sqrt{5} \times \sqrt{5}=5^{1/2} \times 5^{1/2}=5^{1/2+1/2}=5\]

OpenStudy (anonymous):

\[\sqrt{5} * \sqrt{5} = (\sqrt{5})^2\]

OpenStudy (anonymous):

Hint: